# Human-Machine Interfaces for Subsea Telerobotics: From Soda-straw to Natural Language Interactions

### v1

@@ Abstract
This review explores the evolution of human-machine interfaces (HMIs) for subsea telerobotics, tracing back the transition from traditional first-person “soda-straw” consoles (narrow field-of-view camera feed) to advanced interfaces powered by gesture recognition, virtual reality, and natural language models. First, we discuss various forms of subsea telerobotics applications, current state-of-the-art (SOTA) interface systems, and the challenges they face in robust underwater sensing, real-time estimation, and low-latency communication. Through this analysis, we highlight how advanced HMIs facilitate intuitive interactions between human operators and robots to overcome these challenges. A detailed review then categorizes and evaluates the cutting-edge HMI systems based on their offered features from both human perspectives (e.g., enhancing operator control and situational awareness) and machine perspectives (e.g., improving safety, mission accuracy, and task efficiency). Moreover, we examine the literature on bidirectional interaction and intelligent collaboration in terms of sensory feedback and intuitive control mechanisms for both physical and virtual interfaces. The paper concludes by identifying critical challenges, open research questions, and future directions, emphasizing the need for multidisciplinary collaboration in subsea telerobotics.
@@
Subsea telerobotics; marine robotics; human-machine interface; shared autonomy.
@@
Early teleoperated systems were limited in terms of the feedback they provided, often relying on basic visual data from low-resolution cameras. The operator had to rely heavily on experience and intuition to navigate and perform tasks, which increased the cognitive load and risk of errors. While effective for simpler tasks, studies indicate that the limited field-of-view (FOV), appearing as looking through a “soda straw” [14], results in reduced situational awareness in more complex and dynamic environments, leading to increased cognitive load and reduced performance. The 2D visual feedback provides only a partial understanding of the remote workspace, requiring operators to frequently switch between different viewpoints if available [15], and manually integrate information from various perspectives. This challenge is especially pronounced in confined or cluttered spaces, a common scenario in underwater pipelines, caves, and shipwreck exploration. In larger workspaces, traditional monocular camera-based interfaces fail to provide a comprehensive spatial view, leading to significant cognitive overload on the teleoperators.
These limitations have led to the development of more interactive systems including 3D scene rendering with augmented/virtual/mixed reality (XR), haptic cues, and natural language (NL) controls for complex teleoperation tasks. Recent research highlights the importance of human-robot interface design in improving the operator’s awareness, cognitive load, and stress factors [16, 17, 18] – as well as improving the ROVs’ physical capabilities [19]. These developments have allowed operators to engage in human-machine shared autonomy, significantly enhancing the precision and safety of subsea missions [20, 21, 22]. Transparency and reliability are two other critical attributes of shared autonomy, to provide the operator with clear and interpretable insights into the robot’s intentions and actions [23]. Additionally, fostering mutual trust between the operator and the robot is essential, so that the operator feels both in control and supported by the machine’s autonomous capabilities [24].
Fig. 1 presents a general outline of human-machine sensorimotor interaction for subsea telerobotics. As shown, the interface elements bridge the human and machine entities with an interactive flow of teleoperation (teleop) control and feedback commands. Modern interfaces can integrate virtual engines to render a physics-simulated world and digital twin (DT) models of real ROVs [25, 26]. These virtual scenarios, simulated independently [27] or modeled from real-world data [28], are useful for mission training and rehearsal, as real underwater deployments involve significant costs and logistics [29]. DTs also allow for simulating synthetic sensory data and mission logs, supporting the early-stage development and testing of new perception and navigation algorithms in a controlled environment [30, 31, 32].
Modern subsea ROVs integrate various shared autonomy and semi-autonomous control capabilities, where the robot can handle low-level functions such as station-keeping or obstacle avoidance [33, 22]. This blend of human awareness with machine intelligence reduces cognitive workload and allows the operator to focus on high-level tasks [34], improving both operational safety and efficiency. Customizable interfaces that can adapt to the specific mission needs or the skill level of the operator further streamline the process [35], allowing for greater flexibility and adaptability in a wide range of applications.
Figure 1: An outline of the human-robot sensorimotor interaction flow for subsea telerobotics is shown. The interactions occur via control and feedback devices such as visual console, haptic gloves and suit, HMD, as well as by natural languages. A virtual simulation engine is often integrated with the interface to project the remote environment and ROV operation as a digital twin. Simulators with integrated twin models help teleoperation training by mission rehearsal/replay, which is essential for analyzing teleoperators’ cognitive load as well as for prototyping human-machine embodied/shared autonomy features. (FPV: first person view; NL: natural language; XR: augmented, virtual, and mixed reality)
To this end, contemporary studies emphasize the role of feedback realization, cognitive reasoning, and intuitive control modes for mission safety, efficiency, and success [16, 36]. Multimodal feedback has been particularly useful in engaging teleoperators’ sensations through means such as vibrotactile cues on the palm [37], force and tension on the forearm [38], augmented visual displays [39], and textual responses to queries [40]. On the other hand, control commands are typically propagated by traditional buttons, hand gestures [41], body movements [42], or NL instructions [43]. With any/some of these features, a dynamic feedback-control loop significantly increases the sense of telepresence and enhances mission reliability and success rates of complex subsea tasks.
The interdisciplinary nature of modern interface technologies, encompassing haptics, XR, and NL features – has led to a diverse body of research literature. Fig. 2 illustrates a snapshot of subsea HMI technologies developed across academia and industry. These technologies are often built for specific platforms and applications, revealing an absence of a set of standard guidelines. As a result, the body of knowledge remains scattered, with no comprehensive review that compares the technologies, compiles the findings, and highlights opportunities for integration and standardization.
@@
Figure 2: A chronological evolution of subsea HMI technologies is shown. The top row highlights early achievements and milestones. As technology advanced, this century observed industry solutions for interfaces and simulators, incorporating XR and haptic features. In the last five years, NL-based interactions have gained popularity, driven by the success of LLMs.
In this paper, we compile a comprehensive review of the prominent technologies and relevant literature on subsea HMIs over the last few decades. We also identify the knowledge gaps and highlight current trends to facilitate future research. Specifically, we make the following contributions.
  1. 1.
    We categorize the state-of-the-art (SOTA) subsea telerobotic interfaces and simulators based on attributes such as their choice of sensorimotor interfaces, control and feedback mechanisms, platform compatibility, and other relevant features. We then evaluate and compare these features based on their applicability and effectiveness across diverse subsea applications. These discussions are in Section 2-3.
  2. 2.
    We present a detailed review of human-robot interaction and shared autonomy frameworks in subsea telerobotics, highlighting various design issues regarding bidirectional interaction, degree of intelligence, transparency, trust, and the underlying task allocation processes; see Section 4. These analyses also provide valuable insights into human-machine teaming and the associated feedback engineering and interactive control mechanisms; see Section 5.
  3. 3.
    Finally, we analyze the SOTA approaches and current trends to address the unique challenges in subsea HMIs, highlighting the open problems and potential new use cases for future research; see Section 6.
Through this detailed examination of trends, challenges, and potential breakthroughs – we hope to inspire new approaches to bridge the gap between theory and practice. We anticipate that this review will serve as a comprehensive resource for researchers and engineers and support the development of more efficient and reliable telerobotic systems to advance subsea robotics technologies of the next generation.
2 2  Subsea Telerobotic User Interfaces
@@
Telerobotic interfaces have evolved significantly over the past decades, with novel integrated features playing a critical role in enhancing operator control and mission success rate. Traditionally, vision-based interfaces dominated, where operators relied solely on 2D video feeds to navigate and control a robot [4, 44]. Haptic feedback was integrated later [37], improving task precision by allowing operators to feel forces such as drag from fluid dynamics and resistance from interactions with objects. The introduction of XR interfaces [45] further advanced ROV teleoperation by immersing operators in a 3D virtual environment, providing greater situational awareness and enhancing operator control [46, 47]. More recently, researchers have integrated natural language understanding capability into mission management interfaces [48], allowing operators to plan missions and control ROVs using NL commands, making interaction more intuitive.
@@
Table 1: Comparison of existing subsea teleop interfaces based on their integrated features and add-ons. ($\infty$: unlimited number; W: Windows; M: MacOS; L: Linux).
@@
3 2.1   Status Quo of Robot Teleoperation Interface
Subsea HMIs have been developed to address the unique challenges of underwater operations, where visibility, communication bandwidth, localization techniques, and environmental conditions significantly differ from terrestrial systems. The interfaces are referred to as ground control stations (GCS) [61]; examples of which include QGroundControl [49], APM Planner2 [51], EOD Workspace by GreenSea IQ [52], NaviSuite Mobula [56], etc. For smartphones and tablet devices, a few applications are available including QGroundControl, DroidPlanner [60], SidePilot [62], and MAVPilot [63]. However, certain interfaces (e.g., SidePilot and MAVPilot) are tailored exclusively for aerial vehicles, relying on GPS signals that are unavailable underwater. Table 1 compares some key aspects and available features of SOTA telerobotic interfaces.
QGroundControl serves as the official interface for submersible (underwater) vehicles within the ArduPilot flight controller developer community [64]. It is built on ArduSub [65], a comprehensive open-source framework specifically designed for underwater vehicles and integrated into the larger ArduPilot ecosystem. Among proprietary interfaces, leading ROV manufacturer VideoRay offers VideoRay Cockpit [53], exclusively designed to operate their ROVs. NaviSuite Mobula is a third-party interface built for VideoRay ROVs and offers some autonomous capabilities including scanning a wall, orbiting a target, etc. Abyssal Offshore [57] integrates real-time XR visualization, $270\degree270$ ° FOV from multiple cameras, and risk analyses for commercial ROV operation. Open-source platforms such as QGroundControl are more suitable for researchers since they allow customization to suit specific hardware configurations and operational needs. Moreover, OpenROV cockpit [66] offers browser-based connectivity without requiring any software installation. Industry-grade interfaces such as FMC Schilling [67] and Teledyne SeaBotix [54] come as rugged hardware packages and do not require external computers or controllers for teleoperation.
4 2.2   Vision-based Consoles
@@
Underwater vision is a well-explored area in telerobotics literature and has long been the key navigational information for human teleoperators [68]. The environmental challenges as well as the inherent limitations of 2D visual systems have been studied for decades and many innovative solutions have been proposed. We summarize the following challenges and corresponding countermeasures employed in vision-based teleoperation technologies. First, traditional consoles with a fixed POV provide limited information to the operator. Monocular FPV cameras lack depth perception and fail to provide peripheral vision, resulting in significantly reduced situational awareness. Second, low-light turbid water conditions reduce visibility and hinder the capture of clear images. Third, artificial lights used on ROVs to enhance visibility have adverse effects too; their bright light gets reflected and back-scattered by suspended particles directly at the front camera, creating glare and large blind spots for the operator [69].
@@
To compensate for poor visibility and to provide the operator a comprehensive understanding of the scene, modern marine telerobotic systems use high definition wide angle cameras [70], structured lighting systems [71], image enhancement [72], and scene parsing technologies [73]. However, interpreting wide-angle fisheye camera images on 2D displays is difficult because of the distortion effects they introduce. Instead, contemporary works present third person view generation process using augmented reality [14], follower ROV [74], external camera [75], etc. Although not popular yet, multispectral and hyperspectral imaging has the potential to provide more robust vision systems with enhanced marine life detection and mapping mechanisms [76]. While vision will remain the fundamental modality for teleoperation, other sensory augmentations are essential for the immersive realization of the remote environment [13].
Besides underwater ROV’s front camera feed, the visual consoles display a wide range of internal and external sensor readings i.e., telemetry data. External sensors capture information such as pressure, depth, temperature, water current, acoustic scans, etc. Researchers emphasize scanning sonar systems (e.g. side scan sonar, synthetic aperture sonar) for safer navigation, as they detect and display obstacles beyond the limited FOV of cameras [77, 78]. Internal data encompasses the status of onboard electrical and mechanical systems including energy consumption, housing temperature, data recording, and streaming status, as well as the ROV’s pose and velocity. Among these, the attitude and heading reference system (AHRS) estimates and displays the robot’s pose on the console which is crucial for maintaining heading and orientation in feature-deprived underwater scenes [79].
5 2.3   Integration of XR
@@
Virtually experiencing the depths of the ocean provides humans with an immersive experience, enhancing their perception and enabling more effective engagement with remote underwater settings. In $19931993$, NASA launched the first successful Telepresence Remotely Operated Vehicle (TROV) [46] under the Ross Sea ice near the McMurdo Science Station in Antarctica. This project demonstrated the feasibility of streaming video feeds from ROV and rendering computer-generated graphics on an HMD. Preliminary results also showcased ROV control mechanisms via head movements. However, the progress of XR has been slow in the subsea telerobotics domain, unlike its rapid growth in the gaming industry, surgical robotics, and aerial robotics. According to Zhou et al. [80], an XR model should be able to collect, process, transfer, and reconstruct the immersive scene model of the workspace in real-time and enable intuitive robot controls accordingly. The cost and logistics associated with long-term underwater missions and the unique environmental complexities pose challenges to achieving these criteria [81].
@@
Researchers have come up with innovative solutions to address these challenges and have been successful to some extent in XR interfacing between humans and marine robots. DOLPHYN [82], a mixed reality-based aquatic interface for multi-sensory exploration of scuba diving sites, allows a diver to remotely operate an ROV while being aware of own diving trajectory. Multiple contemporary works explore XR-based view synthesis techniques to provide operators with an enhanced peripheral view during teleoperation. For instance, the augmented visualization of the CoMAS project [83] supports ROV pilots during underwater archaeological site exploration and artifacts collection. A similar approach is presented in VENUS project [84], where the archaeological contents are reconstructed in mixed reality by performing bathymetric and photogrammetric surveys on the real site. Moreover, Ego-to-Exo [14] demonstrate the utility of augmented third-person visuals to increase situational awareness during underwater cave exploration; see Fig. 3 (b). In addition to rendering the visuals, an HMD is capable of actively controlling the ROV [39], freeing the operator’s hands for other navigational or manipulation tasks; see Fig. 3 (a). Windowed or stereoscopic VR [46] further improves perception, when compared against monoscopic VR or desktop-based visuals [85]. Other VR-based frameworks such as TWINBOT [86, 87] offer full teleoperation experiences in simulation, which is particularly useful for pilot training prior to real-world missions. The impact of training and XR visuals is supported by studies showing a reduction in task completion time by over 50%, even when the control method remains unchanged [85]. Two key components of XR systems: simulators and digital twins – are further discussed in Sec. 3.
6 2.4   Multimodalities Interface
@@
Besides vision, human sensorimotor control relies on multimodal sensory feedback such as visual, auditory, and somatosensory (tactile and proprioceptive) cues to interpret the consequences of initiated actions [88, 89, 90]. Haptic feedback is the most recognized one in enhancing XR systems; researchers have explored innovative methods for simulating tactile sensations and haptic devices are employed to deliver haptotactile stimulation (e.g., vibrations and force feedback) synchronized with real-time events [91, 92]. Haptic-integrated interfaces typically feature a visual feed via a desktop or a head mounted display (HMD) and one/multiple haptic feedback devices such as wearable gloves, exoskeletons, body suits, etc. Preferred choices for haptic feedback include vibrotactile, skin indent, pressure, etc. on finger and wrist regions as these options only require cheap hardware compared to exoskeletons or full-body haptic suits. Fig. 3 shows a pair of hand-held controllers for vibrotactile feedback. Other haptic feedback mechanisms and their suitable applications are categorized in Table 2.
@@
The benefits of haptic feedback have been extensively studied for ROV driving and manipulation tasks [93, 94]. Ryden et al. [2] present a framework for assisted manipulation (opening and closing of a valve) using real-time haptic feedback to the operator’s hand. It guides the robotic arm to stay away from forbidden regions using force feedback and pulls the arm toward the desired orientation. Such force feedback is crucial for the operator’s sense of virtual presence and is driven by different sources such as artificial muscle [95], magneto-rheological fluid [96, 97], and dc motors [38]. Researchers also explore how other external sensory information can be incorporated into the feedback system. For instance, Amemiya et al. [98] develop a system that combines water pressure and torsion forces to enhance kinesthetic perception. Moreover, the haptic interface developed by Xia et al. [13, 99] includes the feedback of hydrodynamic state as a vector field in addition to other vibrotactile cues, helping the operator to maintain orientation and heading of the ROV in turbulent water.
@@
Furthermore, whole body motion mapping frameworks [42, 100] take advantage of body dexterity and movements to control multiple motions of an ROV. However, as mentioned in Table 2, haptic feedback on multiple sensory organs may reduce user comfort [101]. To address this, researchers include vestibular feedback using an actuated chair that mitigates HMD-induced motion sickness during long-term operations [42]. Additionally, guiding systems are developed for haptic-supported manipulation that evaluate discomfort in human hand using the Rapid Upper Limb Assessment (RULA) [102], and direct the hand toward a more comfortable posture [103]. Overall, despite the limitations of current haptic devices and the low ergonomics in some applications, the combination of visual and haptic feedback have been demonstrated significant potential to enhance the intuitive sense of telepresence.
@@
Table 2: Haptics and XR capabilities of telerobotic interfaces are categorized and their use cases are highlighted.
@@
stylus controlling valve/knob [104, 105]
@@
7 2.5  Integration of Natural Language

### v2

@@ Abstract
This review explores the evolution of human-machine interfaces (HMIs) in subsea telerobotics, charting the progression from traditional first-person “soda-straw" consoles– characterized by narrow field-of-view camera feeds– to contemporary interfaces leveraging gesture recognition, virtual reality, and natural language processing. We systematically analyze the state-of-the-art literature through three interrelated perspectives: operator experience (including immersive feedback, cognitive workload, and ergonomic design), robotic autonomy (contextual understanding and task execution), and the quality of bidirectional communication between human and machine. Emphasis is placed on interface features to highlight persistent limitations in current systems, notably in immersive feedback fidelity, intuitive control mechanisms, and the lack of cross-platform standardization. Additionally, we assess the role of simulators and digital twins as scalable tools for operator training and system prototyping. The review extends beyond classical teleoperation paradigms to examine modern shared autonomy frameworks that facilitate seamless human-robot collaboration. By synthesizing insights from robotics, marine engineering, artificial intelligence, and human factors– this work provides a comprehensive overview of the current landscape and emerging trajectories in subsea HMI development. Finally, we identify key challenges and open research questions and outline a forward-looking roadmap for advancing intelligent and user-centric HMI technologies in subsea telerobotics.
@@
Human-machine interface; subsea telerobotics.
@@
Subsea telerobotic technologies have advanced significantly over the past few decades, driven by the need for remote inspection, maintenance, and research expeditions in underwater environments [1, 2]. Remotely operated vehicles (ROVs) and autonomous underwater vehicles (AUVs) are typically deployed to perform remote tasks in subsea environments that are beyond the reach of human scuba divers [3]. The robots are generally teleoperated from a surface vessel or a base station [4] for applications such as underwater infrastructure inspection [5, 6, 7], seabed mapping [8, 9], remote surveillance [10, 11], environmental monitoring [12], scientific expeditions, and more. Despite growing interest in autonomy, teleoperation continues to dominate subsea missions, comprising over 75% of deployments [13]. This trend highlights the immense potential of advancing telerobotic interfaces for economic growth and scientific discovery.
Early teleoperated systems were limited in terms of the feedback they provided, often relying on basic visual data from low-resolution cameras. The operator had to rely heavily on experience and intuition to navigate and perform tasks, which increased the cognitive load and risk of errors. While effective for simpler tasks, studies indicate that the limited field-of-view (FOV), appearing as looking through a “soda straw” [14], results in reduced situational awareness in more complex and dynamic environments, leading to increased cognitive load and reduced performance. The 2D visual feedback provides only a partial understanding of the remote workspace, requiring operators to frequently switch between different viewpoints if available [15], and manually integrate information from various perspectives. This challenge is especially pronounced in confined or cluttered spaces, a common scenario in underwater pipelines, caves, and shipwreck exploration. In larger workspaces such as coral reefs, offshore rigs, and mining fields– traditional monocular camera-based interfaces fail to provide a comprehensive spatial view, leading to significant cognitive overload on the teleoperators [16, 17].
Figure 1: An outline of the human-robot sensorimotor interaction flow for subsea telerobotics is shown. The interactions occur via control and feedback devices such as visual console, haptic gloves and suit, HMD, as well as by natural languages. A virtual simulation engine is often integrated with the interface to project the remote environment and ROV operation as a digital twin. Simulators with integrated twin models help teleoperation training by mission rehearsal/replay, which is essential for analyzing teleoperators’ cognitive load as well as for prototyping human-machine embodied/shared autonomy features. (FPV: first person view; NL: natural language; XR: augmented/virtual/mixed reality)
These limitations have led to the development of more interactive systems, including 3D scene rendering with augmented/virtual/mixed reality (XR), haptic cues, and natural language (NL) controls for complex teleoperation tasks. Recent research highlights the importance of human-robot interface design in improving the operator’s awareness, cognitive load, and stress factors [18, 19] – as well as improving the ROVs’ physical capabilities [20]. These developments have allowed operators to engage in human-machine shared autonomy, significantly enhancing the precision and safety of subsea missions [21, 22]. Fig. 1 presents a general outline of human-machine sensorimotor interaction for subsea telerobotics. As shown, the interface elements bridge the human and machine entities with an interactive flow of teleoperation (teleop) control and feedback commands. It increases the sense of telepresence via intuitive feedback, integrates intelligent machine capabilities, and offers virtual engines to render a physics-simulated world and digital twin (DT) model of the real ROV [23]. More advanced interfaces offer task-sharing between the operator and the robot, thereby reducing cognitive load and enhancing overall mission efficiency [24].
Knowledge gaps and open questions. The interdisciplinary nature of modern HMIs with haptics, XR, and NL features has led to a broad but fragmented body of research literature. The existing systems and interfaces are built for a diverse set of platforms and applications without any consistent guidelines. There is no comprehensive technical review that compares the interfaces, compiles the findings, and identifies integration opportunities. To this end, we pose the following research questions, highlighting key gaps in HMI design, communication modalities, interaction strategies, and autonomy frameworks.
  1. 1.
    What design principles and standard guidelines are necessary to ensure modularity and cross-platform compatibility of HMIs in subsea telerobotics?
  2. 2.
    What command and feedback modalities (e.g., language, gesture, haptics, camera vision) should be prioritized or fused for different tasks according to complexity, environmental constraints, and mission objectives?
  3. 3.
    To what extent can digital replicas (simulators and twins) mimic a real-world scenario, thereby improving operator skills, reducing mission logistics, and supporting algorithm development for subsea operations?
  4. 4.
    How various aspects of shared autonomy can be integrated into modern interfaces and adapted according to the operator’s skill level and cognitive load?
An in-depth understanding of these aspects is essential for driving innovation and advancing the HMI frontiers of subsea telerobotics. This paper aims to contribute toward that objective by providing a comprehensive technical review of the key HMI dimensions relevant to marine robotic systems.
Specific contributions. In this paper, we present a structured analysis of subsea HMI systems based on their functional roles, interaction modalities, and levels of autonomy as observed in the literature. To facilitate comparative evaluation, we introduce the concept of “technology clusters" that delineate how human-machine interaction happens across different system architectures. First, we categorize and evaluate the state-of-the-art (SOTA) subsea telerobotics interfaces and simulators based on a set of guidelines such as their choice of sensory modalities, platform compatibility, user comfort, and optimal use cases (Section 3-4). Then, we present a detailed review of human-robot teaming in subsea telerobotics, highlighting various aspects of shared autonomy, bidirectional interaction, and intelligence (Section 5-6). Finally, we analyze the SOTA approaches and trends to address the challenges in subsea HMIs, highlighting open problems and use cases for future research (Section 7).
This review is authored from the perspective of researchers actively involved in subsea robotics to bridge the gap between academic advancements and the practical demands of real-world mission scenarios. It aims to provide foundational knowledge to students, researchers, scientists, and engineers engaged in HMI development. The classification and open research questions will offer a roadmap to young researchers for systematic investigation into interface design, autonomy integration, and human-robot collaboration. For industry leaders, the comparative insights into current systems and emerging trends will inform the design of futuristic HMI solutions.
2 2. Review Scope And Criteria
This review spans two primary HMI categories: commercially available interfaces with publicly accessible documentation and peer-reviewed interfaces within the academic literature. Although current efforts toward HMI standardization emphasize the adoption of open-source, widely used, and accessible platforms– such as BlueRobotics hardware [25], Robot Operating System (ROS2) middleware [26], and ArduPilot-based autopilot architectures [27]– these initiatives reflect localized, predominantly U.S.-based community practices. As such, they do not constitute benchmarks that are consistently adopted across the academic and industrial landscapes.
Given the lack of established benchmarks in this domain, we introduce a set of evaluation criteria to systematically assess and compare the diverse range of HMI systems. Specifically, we formulate a set of guidelines grounded in three core aspects: (i) human operators’ cognitive experience with feedback and control modalities, (ii) robot’s degree of autonomy and their relevance for various tasks, and (iii) the quality of bidirectional human-robot communication. Detailed criteria for each evaluation and analysis are presented in the respective sections; itemized features are marked using the following symbols for normalized interpretation:
  • $\dagger$:  Supported by peer review or official documentation
  • $\ddagger$:  Derived from demo/promotional video
@@
  • $\blacklozenge$:  Inferred by the authors based on others’ findings; should be treated as illustrative insights rather than definitive claims.
This approach, while not exhaustive, reflects the broader need for formal standardization efforts (e.g., data formats, open standards, certifications) in HMI technologies.
@@
Figure 2: A chronological evolution of subsea HMI technologies is shown. The top row highlights early achievements and milestones. As technology advanced, this century observed industry solutions for interfaces and simulators, incorporating XR and haptic features (second row). In the last five years, NL-based interactions have gained popularity, driven by the success of LLMs.
3 3. Subsea Telerobotic User Interfaces
@@
Telerobotic interfaces have evolved significantly over the past decades, with novel integrated features playing a critical role in enhancing operator control and mission success rate. Traditionally, vision-based interfaces dominated, where operators relied solely on 2D video feeds to navigate and control a robot [4]. Haptic feedback was integrated later [28], improving task precision by allowing operators to feel forces such as drag and resistance. The introduction of XR interfaces further advanced teleoperation by immersing operators in a 3D virtual environment, providing greater situational awareness [29, 30]. More recently, researchers have integrated natural language understanding capability into interfaces [31], allowing operators to plan missions and control ROVs using NL commands, making interaction more intuitive.
3.1 3.1   Status Quo of Robot Teleoperation Interface
Subsea HMIs have been developed to address the unique challenges of underwater operations, where visibility, communication bandwidth, localization techniques, and environmental conditions significantly differ from terrestrial systems. Fig. 2 organizes key milestones in the evolution of HMI technologies into three distinct eras: (i) foundational space and ocean exploration during the 20th century that laid the base for HMI development, (ii) the emergence of marine robotic companies, novel software and hardware solutions in the early 21st century, and (iii) the recent shift toward immersive VR-integrated and language-driven interfaces. The motivation behind classifying into these eras is not only due to their chronological relevance but also to highlight how technical, commercial, and human factors have co-evolved in shaping the capabilities and expectations of subsea HMI systems.
Generally, an ROV teleoperation interface is referred to as ground control station (GCS) [32]; examples of which include QGroundControl [33], APM Planner2 [34], EOD Workspace by GreenSea IQ [35], NaviSuite Mobula [36], etc. For smartphones and tablet devices, a few applications are available, including QGroundControl, DroidPlanner [37], SidePilot [38], and MAVPilot [39]. However, certain interfaces (e.g., SidePilot and MAVPilot) are tailored exclusively for aerial vehicles, relying on GPS signal that is unavailable underwater. The diverse construction and application domains of these systems make it essential to categorize and evaluate their functionalities from the operator’s perspective. The selected features in Table 1 are chosen to reflect the integration flexibility (e.g., supported OS, open-source availability), built-in safety protocols (e.g., fault diagnostics), and advanced capabilities (e.g., real-time mapping, multi-view 3D displays) for a comprehensive comparison. Note that this table is not exhaustive, particularly for commercial interfaces that do not fully disclose their capabilities. This table is intended to help readers identify the relevant capabilities and make informed decisions when selecting an interface for their use cases.
Among open-source interfaces, QGroundControl is the most popular across the ArduPilot community [40]. It is built on ArduSub [27], a comprehensive open-source framework specifically designed for underwater vehicles. Among proprietary interfaces, leading ROV manufacturer VideoRay offers VideoRay Cockpit [41], exclusively designed to operate their ROVs. NaviSuite Mobula is a third-party interface built for VideoRay ROVs and offers some autonomous capabilities including scanning a wall, orbiting a target, etc. Abyssal Offshore [42] integrates real-time XR visualization, $270\degree270$ ° FOV from multiple cameras, and risk analyses for commercial ROV operation. OpenROV cockpit [43] offers browser-based connectivity without requiring any software installation. Industry-grade interfaces such as FMC Schilling [44] and Teledyne SeaBotix [45] come as rugged hardware packages and do not require external computers or controllers for teleoperation.
@@
Table 1: Comparison of existing subsea teleop interfaces based on their feature and platform compatibility. ($\infty$: unlimited number; W: Windows; M: MacOS; L: Linux); other notations are explained in Sec. 2.
@@
3.2 3.2   Vision-based Consoles
@@
Underwater vision is a well-explored area in telerobotics literature and has long been the key navigational information for teleoperators [50]. The environmental challenges as well as the inherent limitations of 2D visual systems have been studied for decades and many innovative solutions have been proposed. We summarize the following challenges and corresponding countermeasures employed in vision-based interfaces. First, traditional consoles with a fixed monocular POV lack depth perception and fail to provide peripheral vision, resulting in reduced situational awareness. Second, low-light turbid water conditions reduce visibility and hinder the capture of clear images. Third, ROV’s onboard lights get reflected and back-scattered by suspended particles directly at the front camera, creating glare and large blind spots for the operator [51].
@@
To compensate for poor visibility and to provide the operator a comprehensive understanding of the scene, modern marine telerobotic systems use high definition wide angle cameras [52], structured lighting systems [53], image enhancement [54], and scene parsing technologies [55]. However, interpreting wide-angle fisheye camera images on 2D displays is difficult because of the distortion effects they introduce. Instead, contemporary works present third person view generation process using augmented reality [14], follower ROV [56], external camera [57], etc. While vision will remain the fundamental modality for teleoperation, other sensory augmentations are essential for the immersive realization of the remote environment [58].
3.3 3.3   Integration of XR
@@
Virtually experiencing the depths of the ocean provides humans with an immersive experience, enhancing their perception and enabling more effective engagement with remote underwater settings. In $19931993$, NASA launched the first successful Telepresence Remotely Operated Vehicle (TROV) [29] under the Ross Sea ice near the McMurdo Science Station in Antarctica. This project demonstrated the feasibility of streaming video feeds from ROV and rendering computer-generated graphics on a head-mounted display (HMD). Preliminary results also showcased ROV control mechanisms via head movements. However, the progress of XR has been slow in the subsea telerobotics domain, unlike its rapid growth in the gaming industry, surgical robotics, and aerial robotics. According to Zhou et al. [59], an XR model should be able to collect, process, transfer, and reconstruct the immersive scene model of the workspace in real-time and enable intuitive robot controls accordingly. The cost and logistics associated with long-term underwater missions and the unique environmental complexities pose challenges to achieving these criteria.
@@
Researchers have come up with innovative solutions and have been successful to some extent in XR interfacing between humans and marine robots. DOLPHYN [60], a mixed reality-based aquatic interface for multi-sensory exploration of scuba diving sites, allows a diver to remotely operate an ROV while being aware of his own diving trajectory. Contemporary works explore XR-based view synthesis techniques to provide operators with an enhanced peripheral view. For instance, the augmented visualization of the CoMAS project [61] supports ROV pilots during underwater archaeological site exploration and artifacts collection. A similar approach is presented in VENUS project [62], where the archaeological contents are reconstructed in mixed reality by performing bathymetric and photogrammetric surveys on the real site. Moreover, Ego-to-Exo [14] demonstrate the utility of augmented third-person visuals to increase situational awareness during underwater cave exploration; see Fig. 3 (b). In addition to rendering the visuals, an HMD is capable of actively controlling the ROV [63], freeing the operator’s hands for other navigational or manipulation tasks; see Fig. 3 (a). Windowed or stereoscopic VR [29] further improves perception, when compared against monoscopic VR or desktop-based visuals [17]. Other VR-based frameworks such as TWINBOT [64] offer virtual teleoperation, which is particularly useful for pilot training. The impact of training and XR visuals is supported by studies showing a reduction in task completion time by over 50%, even when the control method remains unchanged [17]. Simulators and digital twins for XR systems are further discussed in Sec. 4.
3.4 3.4   Multimodalities Interface
@@
Besides vision, human sensorimotor control relies on multimodal sensory feedback such as auditory and somatosensory (tactile and proprioceptive) cues to interpret the consequences of initiated actions [65]. Haptic feedback is the most recognized one in enhancing XR systems; researchers have explored innovative methods for simulating haptotactile stimulation (e.g., vibrations and force feedback) synchronized with real-time events [66]. Haptic-integrated interfaces typically feature a visual feed via a desktop or an HMD and one/multiple haptic feedback devices such as wearable gloves, exoskeletons, body suits, etc. Preferred choices for haptic feedback include vibrotactile, skin indent, pressure, etc. on finger and wrist regions as these options only require cheap hardware compared to exoskeletons or full-body haptic suits. To clusterize SOTA haptic feedback mechanisms, we define the criteria and terminology used in Table 2 as follows.
  • Feedback location and type represent the anatomical region and sensory modality used to provide physical stimulus to the operator. While multiple locations and feedback types can be integrated in a single system, the classification is based on the dominant feedback point and function.
  • Control mechanism captures the physical console and sensor suites responsible for interpreting and delivering the operator’s command/intention to the machine.
  • Ideal use cases are set by reviewing the user studies and demonstrations reported in literature; metrics such as task relevance and accuracy improvement are considered for this purpose. Note that the examples are not exclusive, rather illustrative of commonly reported subsea applications.
  • Ergonomics reflects a qualitative estimation of the interface’s ease of use and fatigue factors. This is judged based on user studies (where available) and author-reviewed qualitative descriptions from cited works.
@@
The benefits of haptic feedback have been extensively studied for ROV driving and manipulation tasks [67]. Ryden et al. [2] present a framework for assisted valve manipulation using real-time haptic feedback to the operator’s hand. It guides the robotic arm to stay away from forbidden regions and pulls toward the desired orientation using force feedback. To incorporate other external sensory feedback, Amemiya et al. [68] develop a system that combines water pressure and torsion forces to enhance kinesthetic perception. The haptic interface developed by Xia et al. [58, 69] includes the feedback of hydrodynamic state as a vector field, helping the operator to maintain orientation and heading of the ROV in turbulent water.
@@
Whole body motion mapping frameworks [70, 71] take advantage of body dexterity and movements to control multiple motions of an ROV. However, as mentioned in Table 2, haptic feedback on multiple sensory organs may reduce user comfort [72]. To address this, researchers include vestibular feedback using an actuated chair that mitigates HMD-induced motion sickness during long-term operations [70]. Additionally, guiding systems are developed for haptic-supported manipulation that evaluate discomfort in human hand using the Rapid Upper Limb Assessment (RULA) [73], and direct the hand toward a more comfortable posture [74]. Overall, despite current limitations in haptic devices and ergonomics, visuo-haptic feedback has demonstrated significant potential to enhance the intuitive sense of telepresence.
@@
Table 2: Haptics and XR capabilities of telerobotic interfaces are categorized and their use cases are highlighted. Criteria are defined in Sec. 3.4, notations are explained in Sec. 2.
@@
stylus controlling valve/knob [75]
@@
3.5 3.5  Integration of Natural Language